Activity in Limit Theorems



Compute the following limits.



1. lim (4 • f(x))


x→c


2. lim (g(x) - h (x))


x→c ________


3. lim √12 • f(x)


x→c


4. lim (g(x) + h(x)) / f(x)


x→c


5. lim (f(x) + h(x))


x→c




1
Expert's answer
2022-03-10T13:12:13-0500

Solution


Using,


"\mathop {\lim }\limits_{x \to c} f\left( x \right) = \frac{3}{4}\\"


"\mathop {\lim }\limits_{x \to c} g\left( x \right) = 12\\"


"\mathop {\lim }\limits_{x \to c} h\left( x \right) = - 3"



Solution (1)


"\mathop {\lim }\limits_{x \to c} \left[ {4 \cdot f\left( x \right)} \right] = 4\mathop { \cdot \lim }\limits_{x \to c} f\left( x \right) = 4 \cdot \frac{3}{4} = 3\"



Solution (2)


"\mathop {\lim }\limits_{x \to c} \left[ {g\left( x \right) - h\left( x \right)} \right] = \mathop {\lim }\limits_{x \to c} g\left( x \right) - \mathop {\lim }\limits_{x \to c} h\left( x \right) = 12 - \left( { - 3} \right) = 12 + 3 = 15\"



Solution (3)


"\mathop {\lim }\limits_{x \to c} \sqrt {12 \cdot f\left( x \right)} = \sqrt {12 \cdot \mathop {\lim }\limits_{x \to c} f\left( x \right)} = \sqrt {12 \cdot \frac{3}{4}} = \sqrt {3 \cdot 3} = \sqrt 9 = 3\"



Solution (4)


"\mathop {\lim }\limits_{x \to c} \left( {\frac{{g\left( x \right) + h\left( x \right)}}{{f\left( x \right)}}} \right) = \frac{{\mathop {\lim }\limits_{x \to c} g\left( x \right) + \mathop {\lim }\limits_{x \to c} h\left( x \right)}}{{\mathop {\lim }\limits_{x \to c} f\left( x \right)}} = \frac{{12 + \left( { - 3} \right)}}{{\frac{3}{4}}} = 9 \times \frac{4}{3} = 12\"



Solution (5)


"\mathop {\lim }\limits_{x \to c} \left( {f\left( x \right) + h\left( x \right)} \right) = \mathop {\lim }\limits_{x \to c} f\left( x \right) + \mathop {\lim }\limits_{x \to c} h\left( x \right) = \frac{3}{4} + \left( { - 3} \right) = \frac{{3 - 12}}{4} = - \frac{9}{4}\"